OpenStudy (anonymous):

find the lcm of 21y^7 and 63y^3

OpenStudy (anonymous):

21y^7= 3 x 7 x y^7 63y^3= 3 x 3 x 7 x y^3 Therefore LCM = 3 x 3 x 7 x y^7 = 63y^7

OpenStudy (anonymous):

21y^3

OpenStudy (anonymous):

LCM should be number can be divided by those both 2 numbers...